the exterior of a pentagon are in the ratio 1:2:3:4:5. find all the interior angle of the pentagon.

Answer by Alan3354(69443)

the exterior of a pentagon are in the ratio 1:2:3:4:5. find all the interior angle of the pentagon.
-------------
The sum of the exterior angles of all polygons = 360 degrees
---
1x + 2x + 3x + 4x + 5x = 360
15x = 360 degs
x = 24 degs
----------
Each interior angle = 180 - 24*n, n = 1...5
--> 24, 48, 72, 96, 120 degs